OpenStudy (anonymous):

Identify the GCF of the polynomial 6xy - 30x2y4. 6xy 6 6xy - 30x2y4. 5xy

OpenStudy (solomonzelman):

Maybe YOU can take a shot on this one ?

OpenStudy (solomonzelman):

I will give you a hint: \(\normalsize\color{blue}{ 30=\color{red}{6} \times 5 }\) \(\normalsize\color{blue}{ x^2=\color{red}{x} \times x }\) \(\normalsize\color{blue}{ y^4=\color{red}{y} \times y^3 }\)

OpenStudy (solomonzelman):

Maybe YOU can take a shot on this one ?

OpenStudy (anonymous):

I got C when I was working out the problem.

OpenStudy (solomonzelman):

GCF stands for Great Common Factor. The way you answered the question, it seems as though, you are unfamiliar with what GCF means. I am about to leave right now, but maybe someone is willing to give you a lecture about this, if the link I'll give you won't help. http://www.mathsisfun.com/greatest-common-factor.html

OpenStudy (anonymous):

I ws never good at GCF